How do you start solving this equation?y+5/y=-6 ?
Answer by checkley77(12844)

y+5/y=-6 FIRST WE GET RID OF THE FRACTION
(Y*Y+5)/Y=-6
(Y^2+5)/Y=-6 CROSS MULTIPLY
Y^2+5=-6Y
Y^2+6Y+5=0
(Y+5)(Y+1)=0
Y+5=0
Y=-5 ANS.
Y+1=0
Y=-1 ANS.
PROOFS:
-5+5/-5=-5
-5-1=-6
-6=-6
-1+5/-1=-6
-1-5=-6
-6=-6